# modelscsfList of Model Score Tables {CJK*}UTF8gbsn

### v1

@@ Abstract
We tackle the task of long-form music generation—particularly the challenging lyrics-to-song problem—by introducing YuE (乐), a family of open foundation models based on the LLaMA2 architecture. Specifically, YuE scales to trillions of tokens and generates up to five minutes of music while maintaining lyrical alignment, coherent musical structure, and engaging vocal melodies with appropriate accompaniment. It achieves this through: (1) track-decoupled next-token prediction to overcome dense mixture signals, (2) structural progressive conditioning for long-context lyrical alignment, and (3) a multitask, multiphase pre-training recipe to converge and generalize. In addition, we redesign the in-context learning technique for music generation, enabling versatile style transfer (e.g., converting Japanese city pop into an English rap while preserving the original accompaniment) and bidirectional generation. Through extensive evaluation, we demonstrate that YuE matches or even surpasses some of the proprietary systems in musicality and vocal agility. In addition, fine-tuning YuE enables additional controls and enhanced support for tail languages. Furthermore, beyond generation, we show that YuE’s learned representations can perform well on music understanding tasks, where the results of YuE match or exceed state-of-the-art methods on the MARBLE benchmark.
@@ 1 Introduction
Among various music generation tasks, lyrics-to-song audio generation, which involves creating full songs with vocals, accompaniment, from lyrics and control signals, is one of the most challenging. Despite its significance, no open-source system can achieve this at scale. While proprietary systems like Suno, and Udio [^1] https://suno.com/, https://www.udio.com/ have demonstrated impressive results, the lack of open-source alternatives limits accessibility, reproducibility, and innovation. Open-source ecosystems are crucial for advancing AI-driven music generation, enabling collaborative research and laying the foundation for AI models that can understand, compose, and innovate in the arts.
@@ 2 Related Work and Prelimenaries
Audio Tokenizers. Discrete modeling of audio often employs neural codec tokenizers, particularly Residual Vector Quantization GANs (RVQ-GANs) [Kumar et al., 2024], typically categorized into acoustic and semantic tokens [Défossez et al., 2024, Borsos et al., 2023].
Acoustic tokens, optimized for reconstruction, encode fine acoustic details, causing significant token shifts even with minor acoustic variations. Prior studies [Copet et al., 2023b] indicate these tokens require extensive training epochs; notably, we find acoustic tokens alone fail to converge efficiently on our dataset (Section 8.1). Conversely, semantic tokens, derived from self-supervised learning encoders [Schneider et al., 2019, Baevski et al., 2020, Chung et al., 2021, Baevski et al., 2022, Ma et al., 2023], produce semantically meaningful representations (e.g., phonemes, notes, genres) [Zhang et al., 2023, Yuan et al., 2024b, Wang et al., 2025]. Unlike previous work, we conduct extensive experiments on complex in-the-wild music datasets, perform qualitative comparisons, and report tokenizer convergence, demonstrating that fusing semantic information significantly enhance convergence.
@@ 3.1 Overview
Figure 2: Overview of YuE framework: two-stage lyrics-to-song generation with audio/text tokenizers and two language models. Stage-1: music language modeling. Stage-2: residual modeling. Blue: vocal tokens, orange: accompaniment tokens, and grey: residual tokens.
@@ 3.2.1 Track-Decoupled Next-Token Prediction
We quantify Linguistic information Loss After Tokenization (LLAT) using delta word error rate ($\Delta$WER), defined as $\Delta\text{WER}=\text{WER}_{\text{recon}}-\text{WER}_{\text{ori}}$, where $\text{WER}_{\text{recon}}$ and $\text{WER}_{\text{ori}}$ are estimated by a fine-tuned Whisper [^4] A Whisper V3 checkpoint fine-tuned on an internal song dataset with manual transcription. model on tokenizer-reconstructed [^5] We use X-Codec as our tokenizer. See more discussion in Section 3.4 and  8.1. and original mixture audio, respectively. Figure 3 illustrates the relationship between $\Delta$WER and music genre (hip-hop, pop, metal) using 1k sampled tracks. An upward trend is evident, with metal exhibiting the highest LLAT followed by pop and hip-hop, indicating greater modeling difficulty in acoustically dense genres. Vocal-only tracks consistently achieve lower $\Delta$WER compared to mixtures, indicating lower LLAT after source separation.
@@ 3.2.2 Structural Progressive Conditioning
 [^7] We named it “CoT” to pay tribute to the concept of Chain-of-Thought prompting [Wei et al., 2022], as we adopt similar instructions and leverage intermediate conditioning tokens as guidance. However, our approach fundamentally differs from the original Chain-of-Thought prompting in implementation and application. We also acknowledge that there are CoT-like work proposed for audio language models [Du et al., 2024a, Ma et al., 2025, Wang et al., 2025].
To address the long-term decay property issue, we propose an elegant solution that leverages the inherent structural priors of music. Songs are typically composed of distinct segments, such as intro, verse, chorus, bridge, and outro [Nieto et al., 2020, Bruderer et al., 2009, Lerdahl and Jackendoff, 1996]. We use all-in-one [Kim and Nam, 2023] to automatically segment songs into musical sections, with most of the sections shorter than 30 seconds. A song is by average segmented into 14 sessions. Within each structure section, text form segment labels, lyrics and audio are paired together. From a full song perspective, structured text and audio tokens are interleaved (see lyrics2song token arrangement in Figure 4). Special tokens are incorporated to indicate the start and end of the audio.
@@
[…]
@@
Besides, each segment $s_{i}$ is structured as follow:
[…]
@@ 3.2.3 Music In-Context Learning
    Entanglement.  Continuation imposes strong constraints on the style and content of the generated audio. Given music often features structural repetition, the model may simply replicate the reference melody or even entire segments, raising copyright concerns. Moreover, this tight coupling between reference and generated segments diminishes the effectiveness of control prompts or tags designed to steer the creative process.
@@ 3.3 Stage-2: Residual Modeling
As shown in Figure 5, after Stage-1 yields coarse semantic tokens (codebook-0), Stage-2 refines the audio with additional codebooks $1,2,\dots,7$. Denote the total number of codebooks by $K=8$ (indexed from 0 to 7). Although codebook-0 is already produced by Stage-1, we train Stage-2 to predict all codebooks $\{0,1,\dots,7\}$ jointly in a single autoregressive framework. This design ensures that the model has a unified view of both the high-level structure (codebook-0) and the residual details (codebooks 11117777).
Figure 5: The Stage-2 Framework of YuE.
@@ 4.1.1 Multitask Learning
The processed tracks are tokenized and arranged into either tag-conditioned NTP or Dual-NTP formats. Text instructions are prepended before the audio sequences to distinguish the two prediction objectives: Generate music based on the given tags. or Generate music in dual-track format based on the given tags. The tag condition consists of a [genre] string followed by shuffled tags separated by spaces, inserted between the instruction and the audio sequence.
@@ 4.3 Test-time Strategies
Sampling and Classifier-Free Guidance. The sampling parameters are set as follows: top-$k=50$, repetition penalty $=1.1$, top-$p=0.93$, temperature $=1$, and maximum new tokens $=3000$. Classifier-free guidance is applied with a scale of $s=1.5$ for the first segment and $s=1.2$ for subsequent segments  [^14] A lower guidance scale in later segments promotes diversity. to improve the good-case rate. Given the conditional log-probability $\ell_{c}(k)=\log p_{\theta}(k\mid x)$ for token $k$ given prompt $x$ and the unconditional log-probability $\ell_{u}(k)=\log p_{\theta}(k\mid\varnothing)$, the CFG-adjusted log-score is computed as:
@@ 5.1 Data & Training Setup
Training Setup. Our codebase is built upon Megatron-LM [Shoeybi et al., 2019], following the LLaMA2 architecture [Touvron et al., 2023a, Zhang et al., 2024]. Most of our Stage-1 experiments use a 0.5B-scale model trained on 16 NVIDIA H800 GPUs, with a typical token budget of 100B tokens. Under this budget, models usually produce valid outputs, show preliminary lyric-following capabilities, and exhibit basic musical discernment. For scaling experiments, we increase the token budget to 500B and scale models to 0.5B, 2B, and 7B parameters, trained respectively on 32, 96, and 512 NVIDIA H800 GPUs. We maintain a global batch size of 768 when possible by adjusting micro-batch size, gradient accumulation steps, and tensor parallelism; when computational resources are constrained, we reduce the global batch size to 512 or 256. For optimization, we use the Adam optimizer with gradient clipping set to 1.0, weight decay 0.1, $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=10^{-8}$, and parameter initialization with standard deviation 0.02. Detailed training procedures are described in Section 4.1.2.
@@ 6.1 Human Evaluation
To evaluate the subjective musical qualities of YuE and comparative models, we conducted a detailed A/B test on six dimensions: vocal (acoustic) quality, accompaniment (acoustic) quality, music arrangement, melodic attractiveness, vocal-backtrack matching, and song structure. We visualize the win rate with radar plot in Figure 7(L). Suno V4 consistently outperforms all other models across these aspects, thus we normalized the win rate by Suno to improve visual clarity. Among other models, YuE excels notably in music structure and music arrangement, highlighting its capability for coherent long-form composition capability. However, YuE shows clear deficiencies in vocal and accompaniment acoustic quality, likely due to limitations of its current audio tokenization method. While YuE achieves decent musicality and convergence, the semantic-fused tokenizer requires improvements in acoustic detail via, enhanced decoder or super-resolution backend.
@@ 6.2.1 Vocal Agility
As shown in Figure 8, the distribution of song-level vocal ranges across different systems reveals notable variations in vocal agility. Higher values indicate greater vocal expressiveness. Among the models, YuE demonstrates one of the widest vocal ranges (medium =27$\sim$=27∼ = 27 semitones), closely matching top-performing closed-source systems like Suno V4. This suggests that YuE is capable of generating diverse and dynamic vocal performances. In contrast, models like Hailuo and Tiangong show a more constrained vocal range (medium number around 20 semitones), indicating potential limitations in expressiveness. These findings highlight YuE’s strength in producing vocally rich and varied song compositions.
@@ 6.2.3 Model Based Evaluation
Table 3 illustrates model based automatic evaluation results, including distribution metrics KL and FAD, aesthetics metrics proposed by meta, and audio-text alignment score such as CLAP score [Wu et al., 2023a] and CLaMP 3 score [Wu et al., 2025].
Distribution Matching Metrics.
Content Based Metrics.
Alignment Metrics.
YuE attains the highest alignment score according to CLaMP 3 (0.240) [Wu et al., 2025], closely aligning with the human-evaluated c̈ontrolïndicators from the previous section. However, we observe a notably lower alignment for YuE according to the CLAP score (0.118), which not only diverges from human evaluation trends but also directly contradicts the findings from CLaMP 3. These discrepancies highlight potential limitations of the CLAP score in accurately capturing human perceptions of controllability, possibly due to differences in pretraining data and modeling strategies. In contrast, CLaMP 3 appears to benefit from recent methodological improvements and broader, web-scale training resources, resulting in more reliable evaluation outcomes. [^26] We use CLaMP3 as the CLaMP 3 score backend, which is a more recent model compared to CLAP, showing improved results in representation quality and music retrieval tasks due to extensive web-scale pretraining. In contrast, CLAP may suffer from limited exposure to singing/musical content during its training, which could lead to discrepancies in evaluating certain music types.
@@ 6.2.4 Correlation Between Automatic Metrics and Human Evaluation
When considering musicality and average human preference (Table 4), the Vocal Range metric stands out, correlating most strongly (above 0.85) with both subjective ratings. This highlights the central role of vocal expressiveness in shaping listeners’ holistic impressions of generated music. It suggests that the song-level vocal range (or vocal agility) could serve as a relatively reliable proxy for musicality within the lyrics-to-song task, offering a practical measure to capture and quantify vocal expressiveness.
@@ 9 Representation Quality
Table 10: Evaluation of YuE single-track unconditional mode on MARBLE. Including GTZAN genre classification, GS key recognition, MTG top 50 classes tagging, and EMO emotion regression.
@@ 11 Memorization Effect
Following previous literatures [Agostinelli et al., 2023, Yuan et al., 2024a], we investigate whether YuE, in its ICL mode—conditioned on a 30-second audio prompt and original lyrics—reproduces significant portions of its training data. ICL is generally more prone to memorization, making this evaluation critical.
@@
To contextualize these results, we compare them to real-world baselines from GTZAN (genre-level similarities) and Covers80 (known melodic duplicates). Results are shown in Figure 15. The similarity distribution for Ref-Gen pairs is significantly lower than Covers80 and remains moderate even compared to GTZAN. While short repetitive motifs, particularly percussive loops, occasionally occur, overall results indicate that YuE’s ICL mode does not engage in extensive copying. Instead, YuE recombines learned musical patterns creatively, demonstrating that the ICL mode effectively generates original content rather than memorizing entire training samples.

### v2

@@ Abstract
We tackle the task of long-form music generation—particularly the challenging lyrics-to-song problem—by introducing YuE (乐), a family of open foundation models based on the LLaMA2 architecture. Specifically, YuE scales to trillions of tokens and generates up to five minutes of music while maintaining lyrical alignment, coherent musical structure, and engaging vocal melodies with appropriate accompaniment. It achieves this through: (1) track-decoupled next-token prediction to overcome dense mixture signals, (2) structural progressive conditioning for long-context lyrical alignment, and (3) a multitask, multiphase pre-training recipe to converge and generalize. In addition, we redesign the in-context learning technique for music generation, enabling versatile style transfer (e.g., converting Japanese city pop into an English rap while preserving the original accompaniment) and bidirectional generation. Through extensive evaluation, we demonstrate that YuE matches or even surpasses some of the proprietary systems in musicality and vocal agility. In addition, fine-tuning YuE enables additional controls and enhanced support for tail languages. Furthermore, beyond generation, we show that YuE’s learned representations can perform competatively on music understanding tasks, where the results of YuE match or exceed state-of-the-art methods on the MARBLE benchmark.
@@ 1 Introduction
Among various music generation tasks, lyrics-to-song audio generation, which involves creating full songs with vocals, accompaniment, from lyrics and control signals, is one of the most challenging. Despite its significance, no open-source system can achieve this at scale. While proprietary systems like Suno and Udio [^1] https://suno.com/, https://www.udio.com/ have demonstrated impressive results, the lack of open-source alternatives limits accessibility, reproducibility, and innovation. Open-source ecosystems are crucial for advancing AI-driven music generation, enabling collaborative research and laying the foundation for AI models that can understand, compose, and innovate in the arts.
@@ 2 Related Work and Prelimenaries
Audio Tokenizers. Discrete modeling of audio often employs neural codec tokenizers, particularly Residual Vector Quantization GANs (RVQ-GANs) [Kumar et al., 2024], typically categorized into acoustic and semantic tokens [Défossez et al., 2024, Borsos et al., 2023]. Acoustic tokens, optimized for reconstruction, encode fine acoustic details, causing significant token shifts even with minor acoustic variations. Prior studies [Copet et al., 2023b] indicate these tokens require extensive training epochs; notably, we find acoustic tokens alone fail to converge efficiently on our dataset (Section 8.1). Conversely, semantic tokens, derived from self-supervised learning encoders [Schneider et al., 2019, Baevski et al., 2020, Chung et al., 2021, Baevski et al., 2022, Ma et al., 2023], produce semantically meaningful representations (e.g., phonemes, notes, genres) [Zhang et al., 2023, Yuan et al., 2024b, Wang et al., 2025]. Unlike previous work, we conduct extensive experiments on complex in-the-wild music datasets, perform qualitative comparisons, and report tokenizer convergence, demonstrating that fusing semantic information significantly enhance convergence.
@@ 3.1 Overview
Figure 2: Overview of YuE framework: two-stage lyrics-to-song generation with audio/text tokenizers and two language models. Stage-1: music language modeling. Stage-2: residual modeling. Blue: vocal tokens. Orange: accompaniment tokens. Grey: residual tokens.
@@ 3.2.1 Track-Decoupled Next-Token Prediction
[…]
@@ 3.2.2 Structural Progressive Conditioning
 [^7] We named it “CoT” to pay tribute to the concept of Chain-of-Thought prompting [Wei et al., 2022], as we adopt similar instructions and leverage intermediate conditioning tokens as guidance. However, our approach fundamentally differs from the original Chain-of-Thought prompting in implementation and application. We also acknowledge that there is CoT-like work proposed for audio language models [Du et al., 2024a, Ma et al., 2025, Wang et al., 2025].
To address the long-term decay property issue, we propose an elegant solution that leverages the inherent structural priors of music. Songs are typically composed of distinct segments, such as intro, verse, chorus, bridge, and outro [Nieto et al., 2020, Bruderer et al., 2009, Lerdahl and Jackendoff, 1996]. We use all-in-one [Kim and Nam, 2023] to automatically segment songs into musical sections, with most of the sections shorter than 30 seconds. A song is on average segmented into 14 sessions. Within each structure section, text form segment labels, lyrics and audio are paired together. From a full song perspective, structured text and audio tokens are interleaved (see lyrics2song token arrangement in Figure 4). Special tokens are incorporated to indicate the start and end of the audio.
@@
$\texttt{[Genre]\ jazz male deep vocal romantic big band}.$
@@
In addition, each segment $s_{i}$ is structured as follow:
$s_{i}=\textsc{[start\_of\_segment]}\;\circ\;\tau_{i}\;\circ\;\ell_{i}\;\circ\;\textsc{<SOA>}\;\circ\;\psi_{i}\;\circ\;\textsc{<EOA>}\;\circ\;\textsc{[end\_of\_segment]}.$
@@ 3.2.3 Music In-Context Learning
    Entanglement.  Continuation imposes strong constraints on the style and content of the generated audio. Given that music often features structural repetition, the model may simply replicate the reference melody or even entire segments, raising copyright concerns. Moreover, this tight coupling between reference and generated segments diminishes the effectiveness of control prompts or tags designed to steer the creative process.
@@ 3.3 Stage-2: Residual Modeling
As shown in Figure 5, after Stage-1 yields coarse semantic tokens (codebook-0), Stage-2 refines the audio with additional codebooks $1,2,\dots,7$. Denote the total number of codebooks by $K=8$ (indexed from 0 to 7). Although codebook-0 is already produced by Stage-1, we train Stage-2 to predict all codebooks $\{0,1,\dots,7\}$ jointly in a single autoregressive framework. This design ensures that the model has a unified view of both the high-level structure (codebook-0) and the residual details (codebooks 1177).
Figure 5: Stage-2 Framework of YuE. $S$: <SOA>, SS=<SOA>, EE=<EOA>, $S_{i}$=<stage_i>.
@@ 4.1.1 Multitask Learning
The processed tracks are tokenized and arranged into either tag-conditioned NTP or Dual-NTP formats. Text instructions are prepended before the audio sequences to distinguish the two prediction objectives: Generate music based on the given tags or Generate music in dual-track format based on the given tags. The tag condition consists of a [genre] string followed by shuffled tags separated by spaces, inserted between the instruction and the audio sequence.
@@ 4.3 Test-time Strategies
Sampling and Classifier-Free Guidance. The sampling parameters are set as follows: top-$k=50$, repetition penalty =1.1=1.1, top-$p=0.93$, temperature =1=1, and maximum new tokens =3000=3000. Classifier-free guidance is applied with a scale of $s=1.5$ for the first segment and $s=1.2$ for subsequent segments [^14] A lower guidance scale in later segments promotes diversity. to improve the good-case rate. Given the conditional log-probability $\ell_{c}(k)=\log p_{\theta}(k\mid x)$ for token $k$ given prompt $x$ and the unconditional log-probability $\ell_{u}(k)=\log p_{\theta}(k\mid\varnothing)$, the CFG-adjusted log-score is computed as:
@@ 5.1 Data & Training Setup
Training Setup. Our codebase is built upon Megatron-LM [Shoeybi et al., 2019], following the LLaMA2 architecture [Touvron et al., 2023a, Zhang et al., 2024]. Most of our Stage-1 experiments use a 0.5B-scale model trained on 16 NVIDIA H800 GPUs, with a typical token budget of 100B tokens. Under this budget, models usually produce valid outputs, show preliminary lyric-following capabilities, and exhibit basic musical discernment. For scaling experiments, we increase the token budget to 500B and scale models to 0.5B, 2B, and 7B parameters, trained respectively on 32, 96, and 512 NVIDIA H800 GPUs. We further train the 7B LM with additional data, scaling up to a total of 1.75T tokens before starting an annealing phase, during which we apply a 40B-token annealing process. We maintain a global batch size of 768 when possible by adjusting micro-batch size, gradient accumulation steps, and tensor parallelism; when computational resources are constrained, we reduce the global batch size to 512 or 256. For optimization, we use the Adam optimizer with gradient clipping set to 1.0, weight decay 0.1, $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=10^{-8}$, and parameter initialization with standard deviation 0.02. Detailed training procedures are described in Section 4.1.2.
@@ 6.1 Human Evaluation
To evaluate the subjective musical qualities of YuE and comparative models, we conducted a detailed A/B test on six dimensions: vocal (acoustic) quality, accompaniment (acoustic) quality, music arrangement, melodic attractiveness, vocal-backtrack matching, and song structure. We visualize the win rate with radar plot in Figure 7(L). Suno V4 consistently outperforms all other models across these aspects, thus we normalized the win rate by Suno to improve visual clarity. Among other models, YuE excels notably in music structure and music arrangement, highlighting its capability for coherent long-form composition capability. However, YuE shows clear deficiencies in vocal and accompaniment acoustic quality, likely due to limitations of its current audio tokenization method. While YuE achieves decent musicality and convergence, the semantic-fused tokenizer requires improvements in acoustic detail via an enhanced decoder or a super-resolution backend.
@@ 6.2.1 Vocal Agility
As shown in Figure 8, the distribution of song-level vocal ranges across different systems reveals notable variations in vocal agility. Higher values indicate greater vocal expressiveness. Among the models, YuE demonstrates one of the widest vocal ranges (medium $\sim=27$ semitones), closely matching top-performing closed-source systems like Suno V4. This suggests that YuE is capable of generating diverse and dynamic vocal performances. In contrast, models like Hailuo and Tiangong show a more constrained vocal range (medium number around 20 semitones), indicating potential limitations in expressiveness. These findings highlight YuE’s strength in producing vocally rich and varied song compositions.
@@ 6.2.3 Model Based Evaluation
Table 3 illustrates model based automatic evaluation results, including distribution metrics KL and FAD, aesthetics metrics proposed by meta, and audio-text alignment score such as CLAP score [Wu et al., 2023a] and CLaMP 3 score [Wu et al., 2025]. Note that not all metrics align well with human perception. We will further discuss the correlation between each metric and human evaluation in Section 6.2.4.
@@
Distribution Matching Metrics.
Content Based Metrics.
Alignment Metrics.
YuE attains the highest alignment score according to CLaMP 3 (0.240) [Wu et al., 2025], closely aligning with the human-evaluated “control” indicators from the previous section. However, we observe a notably lower alignment for YuE according to the CLAP score (0.118), which not only diverges from human evaluation trends but also directly contradicts the findings from CLaMP 3. These discrepancies highlight potential limitations of the CLAP score in accurately capturing human perceptions of controllability, possibly due to differences in pretraining data and modeling strategies. In contrast, CLaMP 3 appears to benefit from recent methodological improvements and broader, web-scale training resources, resulting in more reliable evaluation outcomes. [^26] We use CLaMP3 as the CLaMP 3 score backend, which is a more recent model compared to CLAP, showing improved results in representation quality and music retrieval tasks due to extensive web-scale pretraining. In contrast, CLAP may suffer from limited exposure to singing/musical content during its training, which could lead to discrepancies in evaluating certain music types.
@@ 6.2.4 Correlation Between Automatic Metrics and Human Evaluation
When considering musicality and average human preference (Table 4), the Vocal Range metric stands out, correlating most strongly (above 0.85) with both subjective ratings. This highlights the crucial role of vocal expressiveness and melodic diversity [^27] One possible explanation relates to AR music generation behavior. Such models often favor high-probability tokens, biasing melodies toward conservative choices like tonic, chord tones, or previously generated notes. Poor optimization (e.g., overfitting) or overly conservative sampling exacerbates this issue, reducing melodic diversity. in listeners’ overall impressions of generated music. We find vocal range to be a practical proxy for musicality and recommend its adoption.
@@ 9 Representation Quality
Table 10: Evaluation of YuE single-track unconditional mode on MARBLE. Including GTZAN genre classification, GS key recognition, MTG top 50 tagging, and EMO emotion regression.
@@ 11 Memorization Effect
Following previous literature [Agostinelli et al., 2023, Yuan et al., 2024a], we investigate whether YuE, in its ICL mode—conditioned on a 30-second audio prompt and original lyrics—reproduces significant portions of its training data. ICL is generally more prone to memorization, making this evaluation critical.
@@
To contextualize these results, we compare them to real-world baselines from GTZAN (genre-level similarities) and Covers80 (known melodic duplicates). Results are shown in Figure 15. The similarity distribution for Ref-Gen pairs is significantly lower than Covers80 and remains moderate even compared to GTZAN. While short repetitive motifs, particularly percussive loops, occasionally occur, overall results indicate that YuE’s ICL mode does not engage in extensive copying. Instead, YuE recombines learned musical patterns creatively, demonstrating that the ICL mode effectively generates original content rather than memorizing training samples.